given the base and sum of the other two sides of triangle,prove that locus of its incentre is an ellipse.

Grade:Upto college level

1 Answers

Ramesh V
14 years ago

Since the base lenght is given and sum of 2 sides is aslo given, lets take the given triangle as follows from an ellipse

where the base is between 2 focii and the other side is P(x,y) a point on ellipse

the 3 points are A(-ae,0)  B(ae,0)   P(x,y)

Lets take given values as                                   base=2ae     [ e is eccentricity of ellipse ]

and                           sum of other 2 sides (PA+PB) =2a 

From properties of ellipse, PB=ePZ  = e(a/e -x) = a-ex

                                                  PA=ePZ  = a+ex  

                                                  AB=2ae

Let I(h,k) be the locus of incentre of triangle ABP.

h = (2aex +(a+ex)ae -ae(a-ex) )/ (2a +2ae)

 h = ex

k = 2aey / 2a(1+e)

k = ey / (1+e)

i.e., x=h/e and y=k(1+e)/e

as we have x2/a2 +y2/b2 =1

                        i.e.,                       h2/a2e2 +k2(1+e)2/a2e2(1-e2) =1

now in the above equation, to prove this is form ELLIPSE   [ NOTE: here I(h,k) is the locus of incentre of triangle ABP.]

put a' =ae

let  (1-e2)/(1+e)2 = 1-e' 2

i.e.,   e' 2= 1 - (1-e2)/(1+e)2 = 2e/(1+e)

now  e' 2= 2e/(1+e) = (e+e)/(1+e) < 1       (as eccentricity of ellipse e<1)

hence e' <1

hence from above we can write in form of ELLIPSE as

h2/(a')2 + k2/(a')2(1-(e')2) =1  which is ellipse

Hence, the problem solved

 

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best Sudarshan.

Regards,

Naga Ramesh

IIT Kgp - 2005 batch